# Opt2Skill: Imitating Dynamically-feasible Whole-Body Trajectories for Versatile Humanoid Loco-Manipulation

### v5

@@ III-B2 Imitation Policy
[…]
@@
human-data $0.04\pm 0.03$ $6.68\pm 8.96$ $0.06\pm 0.08$ $4.25\pm 3.49$ $9.42\pm 8.19$ $4.49\pm 7.35$ $0.11\pm 0.18$ $0.04\pm 0.04$
IK-data $\mathbf{0.02\pm 0.01}$ $7.10\pm 5.73$ $0.06\pm 0.05$ $5.47\pm 2.58$ $5.24\pm 1.54$ $2.09\pm 0.27$ $0.10\pm 0.02$ $0.06\pm 0.01$
Ablation Opt2Skill-hist15151515 $\mathbf{0.03\pm 0.01}$ $5.93\pm 4.81$ $\mathbf{0.05\pm 0.04}$ $3.11\pm 1.58$ $6.48\pm 2.32$ $2.49\pm 1.40$ $0.12\pm 0.06$ $0.05\pm 0.01$
Opt2Skill-hist10101010 $\mathbf{0.03\pm 0.01}$ $\mathbf{4.86\pm 4.18}$ $0.06\pm 0.04$ $\mathbf{2.00\pm 0.76}$ $\mathbf{5.23\pm 1.75}$ $1.14\pm 0.61$ $\mathbf{0.06\pm 0.03}$ $\mathbf{0.02\pm 0.01}$
Opt2Skill-hist5555 $0.04\pm 0.01$ $6.46\pm 4.85$ $0.06\pm 0.05$ $5.63\pm 3.54$ $8.59\pm 4.24$ $\mathbf{1.12\pm 0.96}$ $0.06\pm 0.05$ $0.07\pm 0.01$
Opt2Skill-hist00 $0.04\pm 0.02$ $7.82\pm 5.46$ $0.10\pm 0.09$ $6.68\pm 4.20$ $8.31\pm 4.60$ $1.32\pm 0.66$ $0.07\pm 0.04$ $0.07\pm 0.02$
@@ III-B3 Reward Functions, Domain Randomization, and Curriculum
We design a set of reward functions to guide the robot to track TO-based reference motions and accomplish diverse tasks across various scenarios. The reward functions consist of two main parts: task rewards and regularization rewards, as detailed in Table I. The task rewards guide the humanoid to track commands such as target velocity, end-effector positions, pickup targets, and contact events for loco-manipulation.
Our approach leverages TO to explicitly generate parameterized foot trajectories with controllable step height, length, spacing, and timing, enabling smoother and more robust end-effector tracking and eliminates the need for dense reward shaping or manual design of foot motion heuristics.
@@
To facilitate sim-to-real transfer, we apply domain randomization during training, including randomizing robot dynamics, observation noise, control delays, and external disturbances. The full set of randomized parameters and their sampling ranges is summarized in Table II. We also employ a noise curriculum that gradually increases observation noise and penalty strength. This helps the policy learn more stably in early stages and improves generalization in later stages.
@@ IV-A Robot System
Digit is a general-purpose bipedal humanoid robot developed by Agility Robotics. It weighs approximately $48$ kg and has $30$ degrees of freedom (DoF), including $20$ actuated joints—four in each arm and six in each leg. The passive joints in the ankle and shin are connected via leaf spring four-bar linkages. The toe joints (pitch and roll) are indirectly actuated through rods driven by two motors.
We use a high-fidelity physics-based simulator developed by Agility Robotics to evaluate the trained policies in simulation. The Agility simulator provides realistic dynamics and physical properties of Digit and shares many features with the hardware. Once validated in the Agility simulator, the policies are deployed on hardware.
@@ IV-B Comparison with Different Datasets
  • Human-retargeted motions (Human) Obtained by extracting end-effectors and torso motions from human demonstrations, with joint angles obtained through an inverse kinematics (IK) solve.
  • IK-based trajectories (IK): Generated from hand-designed keypoint sequences (e.g., footstep patterns and torso motion), converted to joint angles via IK.
@@ IV-B1 Dataset Generation
The human-retargeted dataset is selected from the AMASS [41] motion capture dataset, where we choose approximately $70$ trajectories containing non-aggressive forward walking motions. Before applying IK, we calibrate the body shape parameters to align with the morphology of Digit. In addition, we scale the torso speeds to match the new leg length. The torso speed ranges from 00 to $1.5$ m/s, with foot clearance between $10$ and $20$ cm.
To ensure a fair comparison, both IK and TO datasets are generated with matched speed and clearance ranges. The IK dataset is created by specifying target keypoints (feet, hand, torso) and solving IK at each timestep. The TO dataset uses the same target keypoints but solves a whole-body trajectory using DDP. Both the IK and TO datasets contain $2000$ trajectories. All trajectories are generated on flat ground. While the human-retargeted dataset is smaller due to difficulty in generating high-quality retargeted trajectories for the robot, the range of motion parameters is matched across datasets.
@@ IV-B2 Evaluation Metrics
[…]
@@
Figure 3: Success rate comparison across different terrains. Policies trained on human-retargeted, IK-based, and TO-generated trajectories (with matched parameters) are evaluated on varying stair heights (left) and slope angles (right).
@@ IV-B3 Tracking Performance
Table III presents the tracking performance of policies trained with different reference datasets. We evaluate $20$ trajectories per policy. TO and IK trajectories are approximately $14$ seconds long, while human trajectories range from 55556666 seconds. The reported metrics are averaged over the full trajectory duration. Overall, Opt2Skill achieves the best tracking accuracy across most metrics.
Specifically, Opt2Skill achives an average hand tracking error of $2.00$ cm and foot tracking error of $5.23$ cm. In contrast, the human-retargeted policy shows higher errors of $4.25$ cm (hand) and $9.42$ cm (foot), while the IK-based policy yields $5.47$ cm (hand) and $5.24$ cm (foot). Regarding long-term consistency, Opt2Skill achieves the lowest global drift: $1.14\%$ in base position, $0.06$ mm/frame laterally, and $0.02$ mrad/frame in yaw, clearly outperforming the human-retargeted and IK-based policies. Although the IK-based policy achieves the lowest joint angle error ($E_{\text{joint}}=0.02$ rad), its higher task-space tracking errors indicate a fundamental discrepancy between kinematic feasibility and dynamic consistency. These results show that physical feasibility and task adaptability of TO-based references contribute to both local tracking accuracy and long-term global stability.
We further analyze Opt2Skill’s performance under different observation history lengths (00, $5$, $10$, and $15$ steps), as shown in Table III. The results indicate that incorporating a short to moderate observation history significantly improves tracking accuracy, while overly long histories can deteriorate policy performance. Overall, Opt2Skill with a 10-step history achieves the best overall accuracy.
@@ IV-B4 Task Success on Rough Terrain
Fig. 3 shows success rates of the three policies across varying stair heights and slope angles in the Agility simulator. For each policy and each terrain setting, we conduct $10$ trials with difference initial yaw angles uniformly sampled from $-25^{\circ}$ to $25^{\circ}$ to evaluate robustness across diverse orientations ($\theta$ in Fig. 3). Using similar trajectory parameters (base speed $\sim 1$ m/s, foot clearance $\sim 0.2$ m), Opt2Skill consistently outperforms the human-retargeted and IK-based baselines. Specifically, for stair climbing (Fig. 3, left), it maintains high success rates up to $12$ cm steps, while the human and IK-based policies drop sharply beyond $8$ cm. Similarly, for slope traversal (Fig. 3, right), Opt2Skill shows higher robustness at steeper inclinations.
@@ IV-C Evaluating the Role of Torque Information in Contact-Rich Loco-Manipulation
We train all policies using $1400$ TO-generated trajectories with varying desk heights (0.850.850.850.850.950.950.950.95 m) and contact normal forces (0020202020 N).
@@
Force Err. (N) 00 $1.2\pm 2.1$ $\mathbf{0.0\pm 0.0}$ $\mathbf{0.0\pm 0.0}$ $\mathbf{0.0\pm 0.0}$
$5$ $4.1\pm 1.7$ $5.0\pm 0.4$ $3.6\pm 1.7$ $\mathbf{1.1\pm 1.4}$
@@
Top: Contact force tracking error in Newtons. Bottom: Right hand position tracking error in centimeters. Results are averaged over $10$ TO-generated trajectories per group with randomized desk heights (range: 0.850.850.850.850.950.950.950.95 m) for a wiping task.
@@ IV-C1 Performance Analysis
To evaluate the effectiveness of each policy, we conduct six groups of experiments with different reference contact forces levels, ranging from 00 to $25$ N. Each group consists of $10$ TO-generated trajectories with randomized desk heights.
@@
Table. IV shows the contact force and right end effector position tracking errors across varying contact force levels. The Pos policy achieves the best position tracking performance across all scenarios. This is expected, as it optimizes soley for position accuracy without being constrained by contact or torque objectives. However, it produces similar force patterns regardless of the reference force and either fails to generate sufficient contact force (Figs. 4(b)–(f)) or overshoots when no force is required (Fig. 4(a)), due to its lack of contact awareness.
Adding reference contact force to the observation and reward in Pos+F improves average force tracking accuracy, but instability remains in certain conditions. For instance, in Fig. 4(b), Pos+F fails to produce the desired $5$ N of contact force. In contrast, Pos+T, which includes reference torques in the observation and reward, shows improved contact force tracking (average error: $2.7$ N) and better position tracking (average error: $2.9$ cm) compared to Pos+F (Table. IV). While the reference force in Pos+F helps the policy learn how much force to apply, the reference torque in Pos+T provides dynamically grounded information about when and how the force should be applied, resulting in more physically consistent behavior.
Finally, Pos+F+T, which combines both reference contact force and torque signals, achieves the best overall force tracking performance (Table. IV). Its position tracking is also competitive with Pos, and it exhibits the most stable force application across all scenarios (Fig. 4). This highlights the complementary nature of contact force and torque signals, and the critical role of reference torque in enabling generalizable, physically grounded contact behaviors—information uniquely available in TO-generated datasets but absent in IK or human motion datasets.
@@ IV-D Hardware Experiments
For locomotion tasks, Fig. 5 (a) shows accurate base position tracking in the $x$ and $y$ directions during flat ground walking. Fig. 5 (b) and (c) show the foot height as the robot successfully walks up a stair and a ramp, respectively. Notably, our Opt2Skill achieves rough terrain walking capability even though the reference trajectory is designed only for flat ground scenarios. The robot also maintains toe compliance when stepping on stair edges and ramp slopes—enabled by the use of low PD gains in the low-level controller, which also helps reduce impact noise during walking.
@@
Fig. 5 (d) demonstrates a multi-contact whole-body manipulation task, where the robot reaches for an object on a desk located well beyond its support polygon. To achieve this, the robot leans on the table using its elbow to maintain a stable upper body posture, while the opposite arm reaches out to push a box. The legs coordinate with upper-body motion to stabilize and enable the forward-reaching behavior, reflecting the necessity of whole-body motion coordination. The example highlights the ability of our Opt2Skill framework to handle high-dimensional loco-manipulation tasks. In Fig. 5 (e), the Digit robot squats down to pick up a plastic box from a shelf layer (hand height: $0.49$ m) and place it on a higher layer (hand height: $1.76$ m). The plots in Fig. 5 (d) and (e) show the tracking performance of both hand positions during these complex motion sequences. In the desk object pushing task, Opt2Skill achieves an average tracking error of $3.7$ cm for the left hand and $3.8$ cm for the right. In the shelf box pick-up task, the errors are $3.6$ cm for both hands. Notably, a deviation along the $y$-axis occurs when the robot makes contact with the box, caused by a size mismatch between the reference and the real-world box (approximately $2$ cm in the $y$ direction for each hand). All policies are trained using over $1000$ reference trajectories with randomized desk heights (0.850.850.850.850.950.950.950.95 m), hand goals, and motion timings to ensure robustness across task variations and environmental conditions.
@@
Additionally, we conduct four more real-world tasks: pickup+walk, door opening, heavy box pickup, and drawing+wiping. as shown in Fig. 6. In the first task, the robot picks up a box from a standing pose and then walks forward with it. In the second task, the robot pushes open two heavy fire doors using both arms while walking. In the third task, the robot squats down and successfully picks up a box weighing up to $4.9$ kg. Finally, the robot uses its right hand to draw various lines on a whiteboard and then wipes them off using its left hand.
@@
These results illustrate that our framework maintains compliant contact, adapts effectively to environmental variations, and enables the robot to perform everyday loco-manipulation behaviors. Together, these examples further highlight the versatility and generalization capability of the Opt2Skill framework across diverse real-world scenarios.

### v6

@@ III-B2 Imitation Policy
Category Method $E_{\text{joint}}$ (rad) $\downarrow$ $E_{\text{vel-x}}$ (cm/s) $\downarrow$ $E_{\text{vel-yaw}}$ (rad/s) $\downarrow$ $E_{\text{ee-hand}}$ (cm) $\downarrow$ $E_{\text{ee-foot}}$ (cm) $\downarrow$ $E_{\text{pos-x}}$ ($\%$) $\downarrow$ $E_{\text{pos-y}}$ (mm/step) $\downarrow$ $E_{\text{yaw}}$ (mrad/step) $\downarrow$
@@
human-data $0.04\pm 0.03$ $6.68\pm 8.96$ $0.06\pm 0.08$ $4.25\pm 3.49$ $9.42\pm 8.19$ $4.49\pm 7.35$ $11\pm 0.18 0.0$$04\pm 0.04$
IK-data $\mathbf{0.02\pm 0.01}$ $7.10\pm 5.73$ $0.06\pm 0.05$ $5.47\pm 2.58$ $5.24\pm 1.54$ $2.09\pm 0.27$ $10\pm 0.02 0.0$$06\pm 0.01$
Ablation Opt2Skill-hist1515 $\mathbf{0.03\pm 0.01}$ $5.93\pm 4.81$ $\mathbf{0.05\pm 0.04}$ $3.11\pm 1.58$ $6.48\pm 2.32$ $2.49\pm 1.40$ $12\pm 0.06 0.0$$05\pm 0.01$
Opt2Skill-hist1010 $\mathbf{0.03\pm 0.01}$ $\mathbf{4.86\pm 4.18}$ $0.06\pm 0.04$ $\mathbf{2.00\pm 0.76}$ $\mathbf{5.23\pm 1.75}$ $1.14\pm 0.61$ $\mathbf{0.06\pm 0.03}$ $\mathbf{0.02\pm 0.01}$
Opt2Skill-hist55 $0.04\pm 0.01$ $6.46\pm 4.85$ $0.06\pm 0.05$ $5.63\pm 3.54$ $8.59\pm 4.24$ $\mathbf{1.12\pm 0.96}$ $06\pm 0.05 0.0$$07\pm 0.01$
Opt2Skill-hist0 $0.04\pm 0.02$ $7.82\pm 5.46$ $0.10\pm 0.09$ $6.68\pm 4.20$ $8.31\pm 4.60$ $1.32\pm 0.66$ $07\pm 0.04 0.0$$07\pm 0.02$
@@ III-B3 Reward Functions, Domain Randomization, and Curriculum
We design reward functions to guide the robot to track TO-based reference motions and accomplish diverse tasks across various scenarios. The rewards consist of two parts (inspired by  [2, 42]): task and regularization rewards, as detailed in Table I. The task rewards guide the humanoid to track commands such as target velocity, end-effector positions, pickup targets, and contact events for loco-manipulation. Our approach leverages the advantages of the Crocoddyl TO solver to explicitly generate parameterized foot trajectories with controllable step height and timing. These structured, dynamically feasible references improve end-effector tracking and eliminate the need for heuristic reward shaping.
@@
To facilitate sim-to-real transfer, we apply domain randomization during training, including robot dynamics, observation noise, control delays, and external disturbances, inspired by [42] and summarized in Table II. We also employ a noise curriculum that gradually increases observation noise and penalty strength. This helps the policy learn more stably in early stages and improves generalization in later stages.
@@ IV-A Robot System
Digit is a general-purpose bipedal humanoid robot developed by Agility Robotics. It weighs approximately $48$ kg and has $30$ degrees of freedom (DoF), including $20$ actuated joints—four in each arm and six in each leg. Passive joints in the ankle and shin are connected via leaf spring four-bar linkages. The toe joints (pitch and roll) are indirectly actuated through rods driven by two motors. We use a high-fidelity physics-based simulator from Agility Robotics to evaluate the trained policies before hardware deployment. The simulator provides realistic dynamics and physical properties of Digit and shares many features with the hardware.
@@ IV-B Comparison with Different Datasets
  • Human-retargeted motions (Human) Obtained by extracting end-effectors and torso motions from human demonstrations, with joint angles obtained through Quadratic Programming (QP)-based IK.
  • IK-based trajectories (IK): Generated from hand-designed keypoint sequences (e.g., footstep and torso motion), converted to joint angles via QP-based IK.
@@ IV-B1 Dataset Generation
The human-retargeted dataset is selected from the AMASS [43] motion capture dataset, where we choose approximately $70$ trajectories containing non-aggressive forward walking motions. Before applying IK, we calibrate the body shape parameters to align with the morphology of Digit. In addition, we scale the torso speeds to match the new leg length. The torso speed ranges from 0 to $1.5$ m/s, with foot clearance between $10$ and $20$ cm.
To ensure a fair comparison, both IK and TO datasets are generated with matched speed and clearance ranges. The IK dataset is created by specifying target keypoints (feet, hand, torso) and solving IK at each timestep. The TO dataset uses the same target keypoints but solves a whole-body trajectory using DDP. Both the IK and TO datasets contain $2000$ trajectories, all generated on flat ground. While the human-retargeted dataset is smaller due to difficulty in generating high-quality retargeted trajectories for the robot, the range of motion parameters is matched across datasets.
@@ IV-B2 Evaluation Metrics
We evaluate both per-frame tracking and long-term drift. Per-frame metrics include the mean joint angle error $E_{\text{joint}}$ (rad), base linear velocity error $E_{\text{vel-x}}$ (cm/s), and yaw angular velocity error $E_{\text{vel-yaw}}$ (rad/s). We also report hand and foot end-effector tracking errors $E_{\text{ee-hand}}$ and $E_{\text{ee-foot}}$ (cm), which reflect local task-space accuracy. To assess long-term consistency, we report global drift metrics: the normalized final base position drift $E_{\text{pos-x}}$ ($\%$) along the walking direction relative to the reference path length, and the average per-step drift in lateral base position $E_{\text{pos-y}}$ (mm/step) and yaw orientation $E_{\text{yaw}}$ (mrad/step), which reflect accumulated deviations over time.
@@
Figure 3: Success rate comparison across different terrains. Policies trained on human-retargeted, IK-based, and TO-generated trajectories (with matched parameters) are evaluated on stair heights (left) and slope angles (right).
@@ IV-B3 Tracking Performance
Table III presents the tracking performance of policies trained with different reference datasets. We evaluate $20$ trajectories per policy. TO and IK trajectories are approximately $14$ seconds long, while human trajectories range from 5566 seconds. The reported metrics are averaged over the full trajectory duration. Overall, Opt2Skill achieves the best tracking accuracy across most metrics.
Specifically, Opt2Skill achives an average hand tracking error of $2.00$ cm and foot tracking error of $5.23$ cm. In contrast, the human-retargeted policy shows higher errors of $4.25$ cm (hand) and $9.42$ cm (foot), while the IK-based policy yields $5.47$ cm (hand) and $5.24$ cm (foot). Regarding long-term consistency, Opt2Skill achieves the lowest global drift: $1.14\%$ in base position, $0.06$ mm/frame laterally, and $0.02$ mrad/frame in yaw, clearly outperforming the human-retargeted and IK-based policies. Although the IK-based policy achieves the lowest joint angle error ($E_{\text{joint}}=0.02$ rad), its higher task-space tracking errors indicate a fundamental discrepancy between kinematic feasibility and dynamic consistency. These results show that physical feasibility and task adaptability of TO-based references contribute to both local tracking accuracy and long-term global stability. A Wilcoxon signed-rank test confirmed statistically significant differences in end-effector tracking performance between TO and IK/human datasets ($p<0.001$ for both comparisons).
We further analyze Opt2Skill’s performance under different observation history lengths (0, 55, $10$, and $15$ steps), as shown in Table III. The results indicate that incorporating a short to moderate observation history significantly improves tracking accuracy, while overly long histories can deteriorate policy performance. Overall, Opt2Skill with a 10-step history achieves the best overall accuracy.
@@ IV-B4 Task Success on Rough Terrain
Fig. 3 shows success rates of the three policies across varying stair heights and slope angles in the Agility simulator. For each policy and each terrain setting, we conduct $10$ trials with difference initial yaw angles uniformly sampled from $25^{\circ}$ to $25^{\circ}$ to evaluate robustness across diverse orientations ($\theta$ in Fig. 3). Using similar trajectory parameters (base speed $\sim 1$ m/s, foot clearance $\sim 0.2$ m), Opt2Skill consistently outperforms the human-retargeted and IK-based baselines. Specifically, for stair climbing (Fig. 3, left), it maintains high success rates up to $12$ cm steps, while the other policies drop sharply beyond $8$ cm. Similarly, for slope traversal (Fig. 3, right), Opt2Skill shows higher robustness at steeper inclinations.
@@ IV-C Evaluating the Role of Torque Information in Contact-Rich Loco-Manipulation
We train all policies using $1400$ TO-generated trajectories with varying desk heights (0.850.850.950.95 m) and contact normal forces (02020 N).
@@
Force Err.
0 $1.2\pm 2.1$ $\mathbf{0.0\pm 0.0}$ $\mathbf{0.0\pm 0.0}$ $\mathbf{0.0\pm 0.0}$
55 $4.1\pm 1.7$ $5.0\pm 0.4$ $3.6\pm 1.7$ $\mathbf{1.1\pm 1.4}$
@@
Top: Contact force tracking error in Newtons. Bottom: Right hand position tracking error in centimeters. Results are averaged over $10$ TO-generated trajectories per group with randomized desk heights (range: 0.850.850.950.95 m) for a wiping task.
@@ IV-C1 Performance Analysis
To evaluate the effectiveness of each policy, we conduct six groups of experiments with different reference contact forces levels, ranging from 0 to $25$ N. Each group consists of $10$ TO-generated trajectories with randomized desk heights.
@@
Table. IV shows the contact force and right end effector position tracking errors across varying contact force levels. The Pos policy achieves the best position tracking across all scenarios. This is expected, as it optimizes soley for position accuracy without being constrained by contact or torque objectives. However, it produces similar force patterns regardless of the reference force and either fails to generate sufficient contact force (Figs. 4(b)–(f)) or overshoots when no force is required (Fig. 4(a)), due to lack of contact awareness.
Adding reference contact force to the observation and reward in Pos+F improves average force tracking accuracy, but instability remains in certain conditions. For instance, in Fig. 4(b), Pos+F fails to produce the desired $5$ N contact force. In contrast, Pos+T, which includes reference torques in the observation and reward, shows improved contact force tracking (average error: $2.7$ N) and better position tracking (average error: $2.9$ cm) compared to Pos+F (Table. IV). While the reference force in Pos+F helps the policy learn how much force to apply, the reference torque in Pos+T provides dynamically grounded information about when and how to apply it, resulting in more physically consistent behavior.
Finally, Pos+F+T, which combines both reference contact force and torque signals, achieves the best overall force tracking (Table. IV). Its position tracking is also competitive with Pos, and it exhibits the most stable force application across all scenarios (Fig. 4). This highlights the complementary nature of contact force and torque signals, and the critical role of reference torque in enabling generalizable, physically grounded contact behaviors—information uniquely available in TO-generated datasets but absent in IK or human datasets. Moreover, policies trained with TO-generated references adapt when the environment deviates from nominal settings (e.g., table height). For instance, with a trajectory defined for $0.95$ m height and $15$ N contact force, the policy maintained wiping at $0.9$ m and $1.0$ m table heights, achieving average forces of $9.6\pm 1.5$ N and $21.3\pm 2.0$ N, and tracking errors of $4.1\pm 1.6$ cm and $4.5\pm 2.0$ cm, respectively. A Wilcoxon signed-rank test confirmed statistically significant differences in both contact force and hand position tracking errors between Pos+F+T and other baselines across all force levels ($p<0.001$).
The integration of force and torque signals significantly enhances tracking performance, a contribution we believe is essential for high-fidelity imitation of complex dynamic skills. However, we acknowledge this necessitates a more complex pipeline that requires an offline TO stage prior to policy training. We anticipate that future advances in sensing technology and force estimation will simplify the data collection process, further facilitating the adoption of such high-fidelity force signals.
@@ IV-D Hardware Experiments
For locomotion tasks, Fig. 5(a) shows accurate base position tracking in the $x$ and $y$ directions during flat ground walking. We generate the reference trajectories by specifying a fixed gait frequency, step length, and foot clearance, using cubic-spline interpolation to compute desired foot trajectories $\hat{\mathbf{y}}$ described in Sec. III-A. Fig. 5(b) and (c) show the foot height as the robot successfully walks up a stair and a ramp. Notably, Opt2Skill achieves rough terrain walking capability despite the reference being designed only for flat ground scenarios. The robot also maintains toe compliance when stepping on stair edges and ramp slopes—enabled by low PD gains in the low-level controller.
@@
Fig. 5(d) demonstrates a multi-contact whole-body manipulation task, where the robot reaches for an object on a desk located well beyond its support polygon. We generate these reference trajectories via motion phases: (1) one arm moves to the target position; (2) the elbow maintains contact with the desk; (3) the opposite arm reaches the target position and returns; (4) release elbow contact; (5) return torso and hands to the initial pose. These target positions serve as the desired trajectories $\hat{\mathbf{y}}$ in Sec. III-A. To achieve this, the legs coordinate with upper-body motion to stabilize and enable the forward-reaching behavior, reflecting the necessity of whole-body motion coordination. This example highlights the ability of Opt2Skill to handle high-dimensional loco-manipulation tasks. In Fig. 5(e), the robot squats down to pick up a box from a lower shelf and place it on a higher one. Trajectories are generated by: (1) squat down; (2) both arms reach and squeeze to grasp the box; (3) retract arms slightly and stand up; (4) reach to the target and release; (5) return torso and hands to the initial pose. During training, hand targets range from $0.40,\pm 0.10,0.55]~\mathrm{m}$ to $0.50,\pm 0.20,1.70]~\mathrm{m}$. In deployment (Fig. 5(e)), the hands reach $0.49$ m (lower) and $1.76$ m (upper), indicating that the policy can handle moderate variations beyond the training range.
The plots in Fig. 5(d) and (e) show the tracking performance of both hand positions during these complex motion sequences. In the desk object pushing task, Opt2Skill achieves an average tracking error of $3.7$ cm for the left hand and $3.8$ cm for the right. In the shelf box pick-up task, the errors are $3.6$ cm for both hands. Notably, a deviation along the $y$-axis occurs when the robot makes contact with the box, caused by a size mismatch between the reference and the real-world box (approximately $2$ cm in the $y$ direction for each hand).
@@
Additionally, we conduct four more real-world tasks: heavy box pickup, pickup+walk, door opening, and drawing+wiping. as shown in Fig. 6. In the first task, the robot squats down and lifts a box weighing up to $4.9$ kg. the trajectory is generated by predefining grasping goals followed by lifting goals for both hands. In the second task, the robot picks up a box from a standing pose and then walks forward with it, using a trajectory obtained by combining the pickup and locomotion tasks. In the third task, the robot pushes open two heavy fire doors using both arms while walking, with end-effector target positions added to the locomotion trajectory. Finally, the robot uses its right hand to draw various lines on a whiteboard and then wipes them off using its left hand, following a sequence of end-effector subgoals constrained to a planar surface.
@@
These results illustrate that our framework maintains compliant contact, adapts effectively to environmental variations, and enables the robot to perform everyday loco-manipulation behaviors. All policies are trained using over $1000$ reference trajectories per task with randomized parameters, supporting the versatility and generalization capability of the Opt2Skill framework across diverse real-world scenarios.